OpenStudy (anonymous):

What is the derivative of g(x)=(2+(x^2+1)^4)^3?

OpenStudy (anonymous):

\[g(x)=(2+(x^2+1)^4)^3\] You COULD multiply it out and brute force it, but the easy way would be to use chain rule. \[f(x) = g(h(x))\] \[f'(x) = g'(h(x))*h'(x)\] your g() is simply the \[(h(x))^3\] That derivative would yield a \[3*(h(x))^2 * \frac{d(h(x))}{dx}\] The h(x) would be the 2+(x^2+1)^4 which would be another chain rule.

OpenStudy (anonymous):

So, the derivative would be 2((x^2+3)^5+x)(10x(x^2+3)^4+1)?

OpenStudy (anonymous):

remember that \[f(x)=a*x^n\] \[f'(x)=n*a*x^{n-1}\] The power decreases by 1, not increase.

OpenStudy (anonymous):

Wait, uhh...Im kind of confused. let me show you my intermediate steps f'(x)=2((x^2+3)^5 +x) (5(x^2+3)^4+1) =2((x^2+3)^5 +x) (5(x^2+3)^4+1)

OpenStudy (anonymous):

Let me show you the first two steps. I'll call \[x^2+1 = h(x)\] \[g'(x) = 3*(2+h(x)^4)^{3-1}*\frac{d(2+h(x)^4)}{dx}\] \[\frac{d(2+h(x)^4)}{dx}=\frac{d}{dx}2+\frac{d}{dx}h(x)^4=0+4*h(x)^{4-1}*h'(x)=4h(x)^3*h'(x)\]

OpenStudy (anonymous):

Wow, sorry! When I was typing the wrong derivative for the question I asked. :( Haha. but as it follows, 3(2+(x^2+1)^4)^2 (0+4(x^2+1)^3(2x) Right?...

OpenStudy (anonymous):

yes that is correct. I was wondering where you were getting your expression from.

OpenStudy (anonymous):

*I was typing the wrong derivative for the for question. Anyways. Then I would multiply 3 and 4 =12 then12 (2x), which would give me 24x and 24x(x^2+1)^3((x^2+1)^4+2)^2?..

OpenStudy (anonymous):

yes